OpenStudy (anonymous):

Choose an equation for the line that is parallel to the given line and passes through the given point for question 1 and 2. 1. y=5x+10 (2,14) a. y=1/5x+4 b. y=-1/5x-4 c. y=5x-68 d. y= 5x+4 2. y=1/5x-10 (15,-16) a. y= -5x+19 b. y= 1/5x+91/5 c. y=1/5x-19 d. y=5x-19

OpenStudy (anonymous):

@wizfan

OpenStudy (anonymous):

@olivia.smiles @MIS08

OpenStudy (anonymous):

1. D 2. C

OpenStudy (anonymous):

can you help with a few more? and thanks @MIS08

OpenStudy (anonymous):

Tell whether the line for each pair of equations are parallel, perpendicular, or neither for questions 3 and 4. 3. y=-7/8x-1 and -8x+7y=-21 a. parallel b. perpendicular c. neither 4. y=-1/4x+10 and -2x+8y=6 a. parallel b. perpendicular c. neither @MIS08

OpenStudy (anonymous):

Are you giving medals?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

1. the equation is perpendicular 2. Neither

OpenStudy (anonymous):

thanks @MIS08

OpenStudy (anonymous):

Anymore ?

OpenStudy (anonymous):

can you help with english? @MIS08

OpenStudy (anonymous):

Tell me the question and I will see if I can